OpenStudy (desmarie):

A solution has a pH of 2.5. Its OH^- is 3.2 times 10^-12 How is this true?

OpenStudy (desmarie):

@Photon336 @pooja195

OpenStudy (desmarie):

@Whitemonsterbunny17

OpenStudy (photon336):

We must use this relationship. pH + pOH = 14 If we know that the pH = 2.5 the pOH = 11.5 remember p = take the negative log -log of the concentration. if [OH] = 3.2x10^-12 take the -log[3.2x10^-12] and tell me what you get

OpenStudy (photon336):

@desmarie

OpenStudy (desmarie):

I got 11.49485002, which I am assuming rounds up to 11.5.

OpenStudy (desmarie):

@N8KD_Da_Beast

OpenStudy (n8kd_da_beast):

こんにちは(Kon'nichiwa) I'm here!! XD. How can I help u today?

OpenStudy (desmarie):

the question listed above

OpenStudy (n8kd_da_beast):

Well looking at the work u and photon done, I can't help but agree with u.

OpenStudy (desmarie):

but it is because photon is not on and I dont know what to do next

OpenStudy (desmarie):

originally in the question it just gave me the pH. I got the answer to the OH^- by guessing but I want to know how to get the answer correctly

OpenStudy (n8kd_da_beast):

This questions remaining steps have me troubled as well. ごめんなさい(Gomen'nasai) but it seems that I can't be any help at the moment. さようなら(Sayōnara)

OpenStudy (desmarie):

@Abhisar

OpenStudy (abhisar):

Ok, so pH is basically -log[H+]. Right?

OpenStudy (abhisar):

You have been given the concentration of OH- ions. So in order to calculate the pH you first need to calculate the concentration of H+ ions.

OpenStudy (abhisar):

In any solution [H+] x [OH-] = \(\sf 10^{-14} \\ \Rightarrow [H+]=\frac{10^{-14}}{3.2 \times 10^{-12}}=3.12 \times 10^{-3}\)

OpenStudy (abhisar):

Now, pH = -log\((3.12 \times 10^{-3})=2.5\)

OpenStudy (abhisar):

Got it?

OpenStudy (desmarie):

thank you so much

OpenStudy (abhisar):

Welcome c:

OpenStudy (desmarie):

sorry another question. what if it gives me only the pH?

OpenStudy (desmarie):

@Abhisar